Find the extreme values of the following functions.(Write together the problem-solving process.) (1) f(x, y) =...

  1. Find the extreme values of the following functions.(Write together the problem-solving process.)

(1) f(x, y) = x^2 + 2xy + 2y^2

(2) f(x, y) = x^2 + 2xy − 2y^2

(3) f(x, y) = 3x^2 + 6xy − 2y^3
Answer #1

Given, f(x, y) = x + 2xy +242 (0) fx = 2x+2y by = 2x+44 Exy = 2 for fyy To find critical points and ato = 3 +2=0 x4y= zo 2xth(2) f(x4y)= ge²+ 2xy-se? - fx=2xt ry fy = 2x-44 fxy = 2 2 fxx=2 ztrix out fyy = -4 To find critical points :- fu and fy ze =3 f (x,y) = 3x²+ory-243 fx = 6x + by 6x-by² fy fuy = 6 HSL fox = 6 fyy = -124 To find critical points fx=0 and fyzo 6x + 6yoat A (oo) - fyy for = 6 = -124 =-1210) = 0 fyy fny اع fox foly frey (0) (0) -(6)? = 0-36 =-36 (fx. fyy - bez A(0,o) is p sadd